OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). A. y = 7x - 10 B. y = -1/7 x + 2 C. y = 7x - 3 D. y = -7x + 10

OpenStudy (anonymous):

it must A or C because the 7x gives the slope (=7) now you need to plug in x and y values of the point to see if they fit

OpenStudy (anonymous):

wat one do u think itt most likey would be?

OpenStudy (anonymous):

x = 1 where y = -3 try equation A: y = 7x - 10 -3 = 7(1) -10 -3 = -3 so its that one

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

ok - hope you followed the method